The sum of lisa's age and Ellen's age is 41. five years from now, the sum of their ages will be three times lisa's present age. How old is Lisa?
Also to solve this, it has to be either the substitution or elimination method.
Thanks.

Answer by ankor@dixie-net.com(6693)

Try the elimination method
:
Write an equation for each statement:
:
"The sum of lisa's age and Ellen's age is 41."
E + L = 41
:
" five years from now, the sum of their ages will be three times lisa's present age."
(E+5) + (L+5) = 3L
E + L + 10 = 3L
E + L - 3L = -10
E - 2L = -10
Subtract this equation from the 1st equation
E + L = 41
E -2L = -10
--------------subtraction eliminates E, find L
3L = 51
L =
L = 17 yrs is Lisa's present age
:
:
Check solution, E = 41 - 17 = 24
"five years from now, the sum of their ages will be three times lisa's present age."
24 + 5 + 17 + 5 = 3(17)
29 + 22 = 51; confirms our solution